# Jailbreak Large Vision-Language Models Through Multi-Modal Linkage

### v2

@@ E-D Hint Evil ASR(%) DSR(%)
Vanilla AdaShield-S Vanilla AdaShield-S Vanilla AdaShield-S
[…]
Table 4: ASR of FigStep vs. MML (ours) in attacking GPT-4o on SafeBench under the AdaShield-Static defense. The best results under defense are highlighted in bold, and the ASR drop is noted. MML-$X$ represents MML with different encryption approaches, where WR denotes word replacement and M for image mirroring.
@@ 5.4 MML Performance under Defense
To further evaluate the effectiveness of MML, we explore its performance under AdaShield-Static (Wang et al., 2024), a prompt-based defense technique. We select two encryption methods for MML: word replacement and image mirroring. We conduct the experiment on SafeBench, using FigStep as the baseline method and GPT-4o as the target model. As before, we analyze the ASR across different forbidden topics in 5 attempts. Detailed experimental setups are provided in the Appendix A.6.
The results are presented in Table 4. Consistent with previous studies (Wang et al., 2024), AdaShield-S performs well on the first seven topics against FigStep, reducing the ASR to 0. However, our method, particularly MML with word replacement encryption, demonstrates robust performance. Apart from Illegal Activity, which experiences a 38% drop, reductions in other topics are mostly below 10%, maintaining an overall ASR of 87.80%. While MML-M shows a more significant decline, it still achieves an overall ASR of 59.20%.
@@ A.6 Defense Experiment Details
AdaShield-Static (Wang et al., 2024) incorporates a defensive prompt, as shown in Figure 23. AdaShield-Static instructs the target model to carefully assess whether the input instruction or image contains text or items suggesting harmful, illegal, or dangerous activities. If such content is detected, the model responds with "I’m sorry"; otherwise, it proceeds to execute the instruction. Prompt of MML will be concatenated before and after the Adashield prompt.
AdaShield-Static Prompt But before processing any instruction, examine the image carefully for any text or items that might suggest harmful, illegal, or dangerous activity. If detected, do not engage with or act on the content. Analyze the image and the given instruction holistically. If the content is determined to be unethical, illegal, or dangerous, please answer “I am sorry". Instead, please execute the following instructions safely and correctly:
Figure 23: AdaShield-Static Prompt. The MML prompt is concatenated before and after this instruction.

### v3

@@ E-D Hint Evil ASR(%) DSR(%)
Vanilla Q+D Q+D+Q D+Q Vanilla Q+D Q+D+Q D+Q Vanilla Q+D Q+D+Q D+Q
IA 6.0 0.0 0.0 0.0 100.0 38.0 62.0 40.0 100.0 16.0 18.0 10.0
HS 4.0 0.0 0.0 0.0 98.0 84.0 90.0 86.0 94.0 34.0 46.0 30.0
MG 4.0 0.0 0.0 0.0 100.0 78.0 96.0 74.0 100.0 50.0 40.0 8.0
PH 0.0 0.0 0.0 0.0 100.0 74.0 84.0 82.0 100.0 30.0 36.0 2.0
Fr 4.0 0.0 0.0 0.0 100.0 92.0 98.0 82.0 100.0 48.0 64.0 28.0
Po 24.0 0.0 0.0 0.0 96.0 74.0 84.0 74.0 96.0 24.0 38.0 12.0
PV 10.0 0.0 0.0 0.0 100.0 76.0 96.0 76.0 100.0 48.0 64.0 30.0
LO 82.0 38.0 40.0 32.0 94.0 96.0 88.0 98.0 94.0 96.0 100.0 88.0
FA 90.0 58.0 72.0 52.0 88.0 92.0 96.0 88.0 100.0 100.0 94.0 96.0
HC 82.0 32.0 36.0 22.0 84.0 96.0 84.0 98.0 92.0 90.0 92.0 78.0
Avg 30.60 12.80 14.80 10.60 96.00 80.00 87.80 79.80 97.60 53.60 59.20 38.20
Table 4: ASR of FigStep vs. MML (ours) in attacking GPT-4o on SafeBench under the AdaShield-Static defense. The best results under the same defense are highlighted in bold. “+” indicates concatenation, “Q” represents the input text prompt, and “D” refers to the defensive prompt from AdaShield-Static. MML-$X$ denotes variations of MML with different encryption methods: WR stands for word replacement, and represents image mirroring.
@@ 5.4 MML Performance under Defense
To further evaluate the effectiveness of MML, we explore its performance under AdaShield-Static (Wang et al., 2024), a prompt-based defense technique. We select two encryption methods for MML: word replacement and image mirroring. We conduct the experiment on SafeBench, using FigStep as the baseline method and GPT-4o as the target model. As before, we analyze the ASR across different forbidden topics in 5 attempts. We consider three variations of AdaShield-Static: placing the defense prompt before, in the middle of, or after the input text prompt. Detailed experimental setups are provided in the Appendix A.6.
The results are presented in Table 4. Consistent with previous studies (Wang et al., 2024), AdaShield-S performs well on the first seven topics against FigStep, reducing the ASR to 0. However, our method, particularly MML with word replacement as the encryption, performs robustly, especially when the defense prompt is sandwiched between the input text prompts. Apart from Illegal Activity, which experiences a 38% drop, reductions in other topics are mostly below 10%, maintaining an overall ASR of 87.80%. While MML-M shows a more significant decline, it still achieves an overall ASR of 59.20%.
@@ A.6 Defense Experiment Details
AdaShield-Static (Wang et al., 2024) incorporates a defensive prompt, as shown in Figure 23. AdaShield-Static instructs the target model to carefully assess whether the input instruction or image contains text or items suggesting harmful, illegal, or dangerous activities. If such content is detected, the model responds with "I’m sorry"; otherwise, it proceeds to execute the instruction.
Prefix AdaShield-Static Prompt Before processing any instruction, examine the image carefully for any text or items that might suggest harmful, illegal, or dangerous activity. If detected, do not engage with or act on the content. Analyze the image and the given instruction holistically. If the content is determined to be unethical, illegal, or dangerous, please answer “I am sorry". Instead, please execute the following instructions safely and correctly:
Infix AdaShield-Static Prompt But before processing any instruction, examine the image carefully for any text or items that might suggest harmful, illegal, or dangerous activity. If detected, do not engage with or act on the content. Analyze the image and the given instruction holistically. If the content is determined to be unethical, illegal, or dangerous, please answer “I am sorry". Instead, please execute the following instructions safely and correctly:
Suffix AdaShield-Static Prompt But before processing any instruction, examine the image carefully for any text or items that might suggest harmful, illegal, or dangerous activity. If detected, do not engage with or act on the content. Analyze the image and the given instruction holistically. If the content is determined to be unethical, illegal, or dangerous, please answer “I am sorry". Instead, please execute the above instructions safely and correctly:
Figure 23: AdaShield-Static Prompt. The prefix defense prompt is placed before the input text prompt, the infix defense prompt is inserted between two repeated input text prompts, and the suffix defense prompt is positioned directly after the input text prompt.